# DeTurb: Atmospheric Turbulence Mitigation with Deformable 3D Convolutions and 3D Swin Transformers

### v1

@@ 1 Introduction
In this work, we propose a novel framework called DeTurb that integrates two modules: (i) a non-rigid registration module to reduce wavy effects and temporal distortion caused by atmospheric turbulence, and (ii) a feature fusion module to select and fuse useful features for enhanced visualization. Several existing video restoration methods include alignment and feature fusion modules (Wang et al. (2019); Jiang et al. (2023); Lin et al. (2024a, b)), which we found to be more effective when separated rather than combined into a single process. The two-step approach has also proven effective in dealing with atmospheric turbulence issues. Model-based methods, such as space-invariant deconvolution (SID) (Zhu and Milanfar (2013)) and complex wavelet-based fusion (CLEAR) (Anantrasirichai et al. (2013)), employ non-rigid image registration to reduce random perturbations. SID then applies deblurring to the near-diffraction-limited image, while CLEAR produces a sharp image through wavelet-based image fusion. With the advent of deep learning, similar strategies have continued (sometimes referred to as tilt-blur models), such as the CNN-based method AT-Net (Yasarla and Patel (2021)) and the transformer-based method TMT (Zhang et al. (2024)).
In our DeTurb framework, both modules employ UNet-like architectures. The first module learns different levels of turbulence distortion similar to TMT, while the second module extracts features from different scales, reconstructing local details related to semantic meanings. To deal with displacement among frames due to turbulence and moving objects, deformable convolutions have been applied in (Hu and Anantrasirichai (2023)). However, our architecture is distinct in that it processes data in a 3D manner, handling features in both spatial and temporal dimensions simultaneously. Specifically, non-rigid registration is achieved using deformable 3D convolutions (Ying et al. (2020)). After the local spatial alignment among frames, the aligned features from each layer of the pyramid are then enhanced using 3D Swin Transformers (Yang et al. (2023)).
Our main contributions can be summarised as follows:
@@ 2 Existing methods
More recently, transformer-based methods have shown promising results in handling atmospheric turbulence by employing different attention mechanisms to process image patches. This enables precise correction of atmospheric distortions across various scales within an image. These methods have discovered long-term dependencies in data and demonstrated great potential in this field (Mao et al. (2022); Zhang et al. (2023a, 2024, b)). TurbNet (Mao et al. (2022)) extracts features through a transformer UNet-like architecture and utilizes physics-inspired downstream methods to reconstruct clean images. The Swin transformer is employed for estimating turbulent flow in (Zhang et al. (2023a)). TMT (Zhang et al. (2024)) utilizes vision transformers to remove blur, while ASF-Transformer (Zhang et al. (2023b)) integrates spatial-aware and frequency-aware transformer blocks into a UNet framework.
Diffusion models (DMs) are also of interest in these applications. Two recent methods, (Nair et al. (2023)) and (Suin et al. (2024)), focus on restoring faces degraded by atmospheric turbulence using the Denoising Diffusion Probabilistic Model (DDPM), initially proposed by Ho et al. (2020). In (Jaiswal et al. (2023)), a physics-based simulator is directly integrated into the training process of a restoration model. Furthermore, Wang et al. (2023) proposes a conditional diffusion model under a variational inference framework for more generic images. DMs serve as generative priors for blind restoration in (Chung et al. (2023)), which exploits a degradation model involving tilt and blur, akin to the concept in TMT (Zhang et al. (2024)) that utilizes two concatenated modules for tilt removal and deblurring.
@@ 3.1 Non-rigid registration module
As mentioned earlier, several methods address multiple frame inputs of the video using frame or feature alignment processes. Among these, we were inspired by (Zhang et al. (2024)), where a UNet architecture with depth-wise 3D convolutions is used to estimate flow across multiple frames. We improve this process with deformable 3D convolutions. A key contribution is that, in atmospheric turbulent environments, objects exhibit visual distortions within small ranges of pixel displacement, appearing randomly in all directions. The use of deformable 3D convolutions provides flexibility in capturing the shapes of the distorted objects across different scales of the UNet in space and time, enabling the extraction of appropriate features from the distorted scenes. After obtaining multi-scale flows each frame in the input group is warped to the current frame from coarse to fine displacements, as shown in Fig. 1.
@@ 3.3 Loss Functions
Two loss functions are employed similar to (Zhang et al. (2024)): Charbonnier Loss and Edge Loss. In the distortion mitigating module, atmospheric turbulence effects can create outliers in the pixel-wise loss. To address this, Charbonnier loss is used, as it combines the benefits of both $\ell_{1}$ and $\ell_{2}$ losses, effectively handling outliers better (Charbonnier et al. (1997)). Defined by Eq. 2,
@@ 5.1 Experiment settings
We evaluated and compared our method with four state-of-the-art models designed to mitigate atmospheric turbulence in long-range imaging: AT-Net (Yasarla and Patel (2021)), TurbNet (Mao et al. (2022)), TSRWGAN (Jin et al. (2021)), and TMT (Zhang et al. (2024)). Additionally, we included the state-of-the-art video restoration models, BasicVSR++ (Chan et al. (2022)), STA-SUNet (Lin et al. (2024b)), and VRT (Liang et al. (2024)). These models were retrained with our synthetic datasets and tested on both synthetic and real-world datasets. All comparative models were trained using similar data augmentation strategies as employed for the TMT, ensuring comparable conditions and fair performance evaluation.
@@ 5.3 Performance on real dataset
Table 3 shows the average results of these scenes. Comparing the two existing methods that performed best on the synthetic data, BasicVSR++ and TMT, our method achieves better scores across all metrics, particularly in LPIPS. For the no-reference metric NIQE, our method achieves approximately 9% and 16% higher scores, respectively. Fig. 3 illustrates examples of the restored scenes, with the last column showing the pseudo ground truth from CLEAR. Our method produces results with clearer edges and more readable text compared to BasicVSR++ and TMT. Although it cannot restore textures or text as well as CLEAR, our method achieves smoother edges than CLEAR. Overall, DeTurb outperforms other learning-based methods.
@@ 5.4 Ablation study
The results of the ablation study are shown in Table 4. First, we investigated the influence of deformable 3D convolutions by replacing them with depth-wise 3D convolutions, as used in TMT (Zhang et al. (2024)). All metrics indicated a performance reduction, particularly for LPIPS. Next, we tested the impact of the 3D Swin transformation by replacing it with 2D Swin transformers (Liu et al. (2021)). Although the overall performance was reduced, it was not as significant as the decrease observed when replacing deformable 3D convolutions with depth-wise 3D convolutions. This highlights the importance of addressing spatiotemporal distortions caused by atmospheric turbulence. This significance is further confirmed by the substantial reduction in restoration quality when the non-rigid registration module is removed, which has a greater impact than removing the feature fusion module from the pipeline.

### v2

@@ 1 Introduction
In this work, we propose a novel framework called DeTurb that integrates two modules: (i) a non-rigid registration module to reduce wavy effects and temporal distortion caused by atmospheric turbulence, and (ii) a feature fusion module to select and fuse useful features for enhanced visualization. Several existing video restoration methods include alignment and feature fusion modules (Wang et al. (2019); Jiang et al. (2023); Lin et al. (2024a, b)), which we found to be more effective when separated rather than combined into a single process. The two-step approach has also proven effective in dealing with atmospheric turbulence issues. Model-based methods, such as space-invariant deconvolution (SID) (Zhu and Milanfar (2013)) and complex wavelet-based fusion (CLEAR) (Anantrasirichai et al. (2013)), employ non-rigid image registration to reduce random perturbations. SID then applies deblurring to the near-diffraction-limited image, while CLEAR produces a sharp image through wavelet-based image fusion. With the advent of deep learning, similar strategies have continued (sometimes referred to as tilt-blur models), such as the CNN-based method AT-Net (Yasarla and Patel (2021)) and the transformer-based method TMT (Zhang et al. (2024b)).
In our DeTurb framework, both modules employ UNet-like architectures. The first module learns different levels of turbulence distortion, while the second module extracts features from different scales, reconstructing local details related to semantic meanings. A multi-scale approach (like UNet) is essential because the levels of turbulence distortion are unpredictable (atmospheric turbulence is quasi-periodic Li (2009)). Additionally, distortion levels can vary spatially due to the varying distances between objects and the camera—the further an object is to the camera, the more distortion it shows. To address displacement among frames caused by turbulence and moving objects, deformable convolutions were applied in (Hu and Anantrasirichai (2023)). However, our architecture differs by processing data in a 3D manner, handling features in both spatial and temporal dimensions simultaneously. Fluctuations in velocity due to atmospheric turbulence cause local displacements between frames, rather than uniform global shifts (Huebner (2009)). As a result, 3D operations are more effective than 2D ones in extracting and enhancing features. Specifically, non-rigid registration is achieved using deformable 3D convolutions (Ying et al. (2020)). After local spatial alignment among frames, the aligned features from each layer of the pyramid are enhanced using 3D Swin Transformers (Yang et al. (2023)).
In summary, our main contributions can be summarised as follows:
@@ 2 Existing methods
More recently, transformer-based methods have shown promising results in handling atmospheric turbulence by employing different attention mechanisms to process image patches. This enables precise correction of atmospheric distortions across various scales within an image. These methods have discovered long-term dependencies in data and demonstrated great potential in this field (Mao et al. (2022); Zhang et al. (2023a, 2024b, b)). TurbNet (Mao et al. (2022)) extracts features through a transformer UNet-like architecture and utilizes physics-inspired downstream methods to reconstruct clean images. The Swin transformer is employed for estimating turbulent flow in (Zhang et al. (2023a)). TMT (Zhang et al. (2024b)) utilizes vision transformers to remove blur, while ASF-Transformer (Zhang et al. (2023b)) integrates spatial-aware and frequency-aware transformer blocks into a UNet framework.
Diffusion models (DMs) are also of interest in these applications. Two recent methods, (Nair et al. (2023)) and (Suin et al. (2024)), focus on restoring faces degraded by atmospheric turbulence using the Denoising Diffusion Probabilistic Model (DDPM), initially proposed by Ho et al. (2020). In (Jaiswal et al. (2023)), a physics-based simulator is directly integrated into the training process of a restoration model. Furthermore, Wang et al. (2023) proposes a conditional diffusion model under a variational inference framework for more generic images. DMs serve as generative priors for blind restoration in (Chung et al. (2023)), which exploits a degradation model involving tilt and blur, akin to the concept in TMT (Zhang et al. (2024b)) that utilizes two concatenated modules for tilt removal and deblurring.
@@ 3.1 Non-rigid registration module
As mentioned earlier, several methods address multiple frame inputs of the video using frame or feature alignment processes. Among these, we were inspired by (Zhang et al. (2024b)), where a UNet architecture with depth-wise 3D convolutions is used to estimate flow across multiple frames. We improve this process with deformable 3D convolutions. A key contribution is that, in atmospheric turbulent environments, objects exhibit visual distortions within small ranges of pixel displacement, appearing randomly in all directions. The use of deformable 3D convolutions provides flexibility in capturing the shapes of the distorted objects across different scales of the UNet in space and time, enabling the extraction of appropriate features from the distorted scenes. After obtaining multi-scale flows each frame in the input group is warped to the current frame from coarse to fine displacements, as shown in Fig. 1.
@@ 3.3 Loss Functions
Two loss functions are employed similar to (Zhang et al. (2024b)): Charbonnier Loss and Edge Loss. In the distortion mitigating module, atmospheric turbulence effects can create outliers in the pixel-wise loss. To address this, Charbonnier loss is used, as it combines the benefits of both $\ell_{1}$ and $\ell_{2}$ losses, effectively handling outliers better (Charbonnier et al. (1997)). Defined by Eq. 2,
@@ 5.1 Experiment settings
We evaluated and compared our method with four state-of-the-art models designed to mitigate atmospheric turbulence in long-range imaging: AT-Net (Yasarla and Patel (2021)), TurbNet (Mao et al. (2022)), TSRWGAN (Jin et al. (2021)), TMT (Zhang et al. (2024b)), and DATUM (Zhang et al. (2024a)). Additionally, we included the state-of-the-art video restoration models, BasicVSR++ (Chan et al. (2022)), STA-SUNet (Lin et al. (2024b)), and VRT (Liang et al. (2024)). These models were retrained with our synthetic datasets and tested on both synthetic and real-world datasets. All comparative models were trained using similar data augmentation strategies as employed for the TMT, ensuring comparable conditions and fair performance evaluation.
@@ 5.3 Performance on real dataset
Table 3 shows the average results of these scenes. Comparing the two existing methods that performed best on the synthetic data, BasicVSR++ and TMT, our method achieves better scores across all metrics, particularly in LPIPS. For the no-reference metric NIQE, our method achieves approximately 9% and 16% higher scores, respectively. Fig. 3 and Fig. 4 illustrate examples of the restored scenes, with the last column showing the pseudo ground truth from CLEAR. Our method produces results with clearer edges and more readable text compared to BasicVSR++ and TMT. Although it cannot restore textures or text as well as CLEAR, our method achieves smoother edges than CLEAR. Overall, DeTurb outperforms other learning-based methods.
@@
Figure 4: Subjective results (similar to Fig. 11 in the DATUM paper).
@@ 5.4 Ablation study
The results of the ablation study are shown in Table 4. First, we investigated the influence of deformable 3D convolutions by replacing them with depth-wise 3D convolutions, as used in TMT (Zhang et al. (2024b)). All metrics indicated a performance reduction, particularly for LPIPS. Next, we tested the impact of the 3D Swin transformation by replacing it with 2D Swin transformers (Liu et al. (2021)). Although the overall performance was reduced, it was not as significant as the decrease observed when replacing deformable 3D convolutions with depth-wise 3D convolutions. This highlights the importance of addressing spatiotemporal distortions caused by atmospheric turbulence. This significance is further confirmed by the substantial reduction in restoration quality when the non-rigid registration module is removed, which has a greater impact than removing the feature fusion module from the pipeline.